INCOME STATEMENT Byron Books Inc. recently reported $12 million of net income. Its EBIT was $26.4 million, and its tax rate was 40%. What was its interest expense? [Hint: Write out the headings for an income statement, and then fill in the knowrn values. Then divide $12 million of net income by (1-T) 0.6 to find the pretax income. The difference between EBIT and taxable income must be interest expense. Use this same procedure to complete similar problems.] Write out your answer completely. For example, 25 million should be entered as 25,000,000. Round your answer to the nearest dollar, if necessary. Do not round intermediate calculations
Answer #1

The interest expense can be calculated as :

As per the income statement :

EBIT

less interest

EBT

Less taxes

NET INCOME

Let the interest expense be X,

(26.4 - X) *0.6 = $12

solving for x we get,

X = $6.4 million which is $6400,000.

Or, EBIT - taxable income is interest

taxable income is= net income/ taxes '

= 12/0.6 = $20

so 26.4 - 20 = $6.4million or 64,000,00 is the interest expense.